OpenStudy (anonymous):

find two consecutive integers whose sum is 67. Which of the following equations could be used to solve the problem? 2x = 67 2x + 1 = 67 2x + 2 = 67 x2 + 1= 67

OpenStudy (anonymous):

Let's call the unknown number 'x'. We know for 2 numbers to be consecutive, one has to be one more than the other (i.e. 4, and 5 are consecutive numbers, where 5 is 1 greater than 4). Therefore: x + (x+1) = 67 ^ ^ Integer 2 Integer 1 By simplifying, we get 2x + 1 = 67. Therefore, either "2x + 1 = 67" or "x2 + 1= 67" can be used. Your answer by solving for 'x' would be: 2x=66 x=33 Therefore, (x+1)=34 Hence, the 2 consecutive integers are 33 and 34. When added together, they equal 67, and they are consecutive (34 is 1 larger than 33).

OpenStudy (anonymous):

so would it be a. b. c. or d.

OpenStudy (anonymous):

d i guess...

OpenStudy (anonymous):

Yeah Or B both get you to the same spot